# Direct optimization of neoclassical ion transport in stellarator reactors

### v1

@@ 4 Conclusions
We have shown through several examples that neoclassical electron and fuel ion transport can be optimized separately to raise the ratio $L_{11}^{e}/L_{11}^{i}$ and produce an outward-pointing ambipolar electric field in the core under reactor-relevant conditions. These configurations are expected to expel heavy impurities from the core during steady-state operation. However, we were unable to generate any completely satisfactory configurations despite substantial effort. One configuration we produced had stronger fuel-ion-relevant $D_{11}^{*}$ suppression than any of the configurations presented here, but effects related to the slope of the $D_{11}^{*}$ vs $\nu^{*}$ curves (like those in figure 2) and the resonant electric field prevented an electron root from being realized. Further development is needed before the approach outlined in this work is ready for mainstream use.
A possible extension of this work involves coupling a modern, fast neoclassical transport code such as KNOSOS [57, 58] to a modern optimization framework such as SIMSOPT [59] and performing optimizations similar to those presented in section 3. This will likely avoid some of the numerical difficulties we encountered while attempting to improve the optimizations and allow for consideration of additional neoclassical effects. Notably, KNOSOS includes the tangential magnetic drifts (unlike DKES), which can be substantial for some configurations [60] and alter the confinement of fuel and impurity ions [61]. The tangential drifts can also make the electron root more difficult to achieve in maximum-$J$ configurations [62]. Because the maximum-$J$ property helps mitigate turbulence [63], it may be a useful optimization target for future devices. A code that takes the tangential drifts into account would therefore be helpful if an electron root is to be targeted as well. Due to its speed, KNOSOS can also calculate thermal transport coefficients very quickly. The final stage of the optimization in section 3.1 successfully improved a configuration with optimized monoenergetic transport coefficients by targeting thermal transport coefficients using the (relatively slow) SFINCS code. This suggests that targeting thermal transport coefficients throughout the optimization may be helpful — Lascas Neto and coworkers [64] have recently used SFINCS to do so and obtained promising results.
Another possible extension of this work involves developing a more theoretical approach to ensuring excellent neoclassical confinement of fuel ions and moderate confinement of electrons. Helander and coworkers [28] exploited the fact that in the formulation of omnigenity, there are separate conditions for zero average radial drift of marginally trapped particles (which pertains to fuel ions) and deeply trapped particles (which pertains to electrons). They derive an expression that allows the former conditions to be satisfied while the latter are violated and optimize a stellarator configuration using this expression as an objective. We suggest that further refinement of the analytical proxy, or perhaps a blend of the analytical and computational approaches, may provide the most reliable results for future stellarator optimization efforts.
@@ Acknowledgments
We thank Stefan Buller, José Luis Velasco, Matt Landreman, Per Helander, Michael Drevlak, Alan Goodman, Eduardo Lascas Neto, Eduardo Rodríguez, and Katia Camacho Mata for their insights and assistance.

### v2

@@ 5 Conclusions
We have shown by example that neoclassical electron and fuel ion transport can be optimized separately in stellarators to raise the ratio of thermal transport coefficients $L_{11}^{e}/L_{11}^{i}$ and produce an outward-pointing ambipolar electric field (“electron root”) in the core under reactor-relevant conditions. Configurations with an electron root are expected to expel heavy impurities from the core during steady-state operation. We control $L_{11}^{e}$ through $\epsilon_{\mathrm{eff}}$ and $L_{11}^{i}$ through an array of fuel-ion-relevant monoenergetic transport coefficients ($D_{11}^{*}$); both $\epsilon_{\mathrm{eff}}$ and $D_{11}^{*}$ are characteristic of a given magnetic field and do not depend on plasma profiles. While our results are encouraging, further development is likely needed before the presented algorithm is ready for mainstream use.
@@ Acknowledgments
We thank Stefan Buller, José Luis Velasco, Matt Landreman, Per Helander, Michael Drevlak, Alan Goodman, Eduardo Lascas Neto, Eduardo Rodríguez, and Katia Camacho Mata for their insights and assistance. We also thank the anonymous referees, whose thoughtful feedback substantially improved this work.